How many samples of size n=3 can be drawn from this population?

1
Expert's answer
2022-02-24T11:03:57-0500

The population size is "N=12". The number of possible samples of size "n=3"  that can be drawn from this population without replacement is "\binom{N}{n}=\binom{12}{3}=220"  

The number of samples that can be drawn from this population with replacement is "N^n=12^3=1728"